# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v2

@@ 4 Experiments
$T=2.0$ $260\pm 0.13$ $210\pm 0.43$ $\mathbf{270\pm 0.18}$
$T=2.1$ $250\pm 0.14$ $208\pm 0.41$ $\mathbf{255\pm 0.19}$
$T=2.2$ $239\pm 0.16$ $210\pm 0.39$ $\mathbf{252\pm 0.17}$
$T=2.3$ $231\pm 0.15$ $214\pm 0.37$ $\mathbf{233\pm 0.17}$
$T=2.4$ $\mathbf{225\pm 0.17}$ $212\pm 0.37$ $\mathbf{225\pm 0.15}$
$T=2.5$ $219\pm 0.17$ $202\pm 0.37$ $\mathbf{221\pm 0.14}$
$T=2.6$ $\mathbf{216\pm 0.18}$ $181\pm 0.40$ $214\pm 0.14$
$T=2.7$ $\mathbf{212\pm 0.18}$ $189\pm 0.36$ $\mathbf{212\pm 0.14}$
Ising model Finally, we verify the performance of EDG on the 2-dimensional Ising model [17], a mathematical model of ferromagnetism in statistical mechanics. To ensure the continuity of physical variables, we employ a continuous relaxations trick [53] to transform discrete variables into continuous auxiliary variables with the target distribution:
$\pi(\boldsymbol{x})=\exp\left(-\frac{1}{2}\boldsymbol{x}^{T}\left(K(T)+\alpha I\right)^{-1}\boldsymbol{x}\right)\times\prod_{i=1}^{N}\cosh\left(x_{i}\right),$
where $K$ is an $N\times N$ symmetric matrix depending on the temperature $T$, $\alpha$ is a constant guaranteeing $K+\alpha I$ to be positive. For the corresponding discrete Ising variables $\boldsymbol{s}=\{1,-1\}^{\otimes N}$, one can directly obtain discrete samples according to $\pi(\boldsymbol{s}|\boldsymbol{x})=\prod_{i}\left(1+e^{-2s_{i}x_{i}}\right)^{-1}$. When there is no external magnetic field and each spin can only interact with its neighboring spins, $K$ is defined as $\sum_{<ij>}s_{i}s_{j}/T$, with the nears neighboring sum $<ij>$. Consequently, the normalizing constant of the continuous relaxations system is given by $\log Z=\log Z_{\text{Ising }}+\frac{1}{2}\ln\operatorname{det}(K+\alpha I)-\frac{N}{2}[\ln(2/\pi)-\alpha]$ [17]. Additionally, using the method described in Sec. 3.3, we provide lower bound estimates of $\log Z_{\text{Ising}}$ at different temperatures for samples generated by NeuralRG, PIS and EDG. Since these are lower bound estimates, larger values indicate more accurate results. As seen in Table 4, EDG provides the most accurate estimates of $\log Z$ across most temperature ranges. The generated states at different temperatures are displayed in Fig. 3.
@@ 5 Conclusion
In summary, our work introduces the EDG as an innovative and effective sampling approach by combining principles from VI and diffusion based methods. Drawing inspiration from VAEs, EDG excels in efficiently generating samples from intricate Boltzmann distributions. Leveraging the expressive power of the diffusion model, our method accurately estimates the KL divergence without the need for numerical solutions to ordinary or stochastic differential equations. Empirical experiments validate the superior sampling performance of EDG.
In consideration of its powerful generation ability and unrestrained network design theoretically, there is still room for further exploration. We can design specific network architectures for different tasks and find the most elaborate one. Nevertheless, it is the first attemp to design a one-shot generator assisted by diffusion models.
In future, our primary focus will be on extending the application of EDG to construct generative models for large-scale physical and chemical systems, such as proteins [54, 55].
@@ Acknowledgements
The first and third authors are supported by the NSF of China (under grant number 12171367). The second author is supported by the NSF of China (under grant numbers 92270115, 12071301) and the Shanghai Municipal Science and Technology Commission (No. 20JC1412500), and Henan Academy of Sciences. The last author was supported by the NSF of China (No. 12288201), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDA25010404), the National Key R&D Program of China (2020YFA0712000), the Youth Innovation Promotion Association, CAS, and Henan Academy of Sciences.
Appendix A Proof of the independence between $z_{T}$ and $x$
In the decoding process, if $z_{T}$ is independent of $z_{0}$, we have
$p_{D}(z_{T}|x)$ $=$ $\int p_{D}(z_{T}|z_{0})p_{D}(z_{0}|x)\mathrm{d}z_{0}$
$=$ $p_{D}(z_{T})\cdot\int p_{D}(z_{0}|x)\mathrm{d}z_{0}$
$=$ $p_{D}(z_{T}).$
Appendix B Proof of (5)
For an extremely small lag time $\tau$, the Euler-Maruyama discretization of $z_{[\cdot]}$ in the decoding process the provides
$z_{t}$ $=$ $z_{t-\tau}+f(z_{t-\tau},t-\tau)\tau+\sqrt{\tau}g(t-\tau)u_{t-\tau},$ (9)
$z_{t-\tau}$ $=$ $z_{t}-\left(f(z_{t},t)-g(t)^{2}\nabla\log p_{D}(z_{t})\right)\tau+\sqrt{\tau}g(t)\bar{u}_{t},$ (10)
where $u_{t-\tau},\bar{u}_{t}\sim\mathcal{N}(\cdot|0,I)$. Importantly, $u_{t-\tau}$ is independent of $\{z_{t^{\prime}}|t^{\prime}\leq t-\tau\}$ and $x$, and $\bar{u}_{t}$ is independent of $\{z_{t^{\prime}}|t^{\prime}\geq t\}$. The Euler-Maruyama approximation of $p_{E}(z_{[\cdot]}|x)$ gives
$z_{t-\tau}=z_{t}-\left(f(z_{t},t)-g(t)^{2}s\left(z_{t},x,t\right)\right)\tau+\sqrt{\tau}g(t)v_{t}$ (11)
with $v_{t}\sim\mathcal{N}(\cdot|0,I)$.
According to (10, 11), we can obtain
$\mathbb{E}_{p_{D}}\left[\log p_{D}(z_{t-\tau}|z_{t})\right]$ $=$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}$ (12)
@@ Appendix B Proof of (5)
$-\frac{1}{2}\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left\|z_{t-\tau}-z_{t}+\left(f(z_{t},t)-g(t)^{2}\nabla\log p_{D}(z_{t})\right)\tau\right\|^{2}\right]$
$=$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}-\frac{1}{2}\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left\|\sqrt{\tau}g(t)\bar{u}_{t}\right\|^{2}\right]$
$=$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}-\frac{d}{2}$
and
$\mathbb{E}_{p_{D}}\left[\log p_{E}(z_{t-\tau}|z_{t},x)\right]$ $=$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}$ (13)
$-\frac{1}{2}\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left\|z_{t-\tau}-z_{t}+\left(f(z_{t},t)-g(t)^{2}s\left(z_{t},x\right)\right)\tau\right\|^{2}\right]$
$=$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}$
$-\frac{1}{2}\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left\|z_{t-\tau}-z_{t}+f(z_{t},t)\tau\right\|^{2}\right]$
$-\frac{1}{2}\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left\|g(t)^{2}s\left(z_{t},x\right)\tau\right\|^{2}\right]$
$+\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left(z_{t-\tau}-z_{t}+f(z_{t},t)\tau\right)^{\top}g(t)^{2}s\left(z_{t},x\right)\tau\right]$
$=$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}$
$-\frac{1}{2}\tau^{-1}g(t)^{-2}\mathbb{E}_{p_{D}}\left[\left\|g(t)^{2}\nabla\log p_{D}(z_{t})\tau+\sqrt{\tau}g(t)\bar{u}_{t}\right\|^{2}\right]$
$-\frac{1}{2}\tau g(t)^{2}\mathbb{E}_{p_{D}}\left[\left\|s\left(z_{t},x\right)\right\|^{2}\right]$
$+\mathbb{E}_{p_{D}}\left[\left(z_{t-\tau}-z_{t}+f(z_{t},t)\tau\right)^{\top}s\left(z_{t},x\right)\right]$
=$\overset{(i)}{=}$ $-\frac{d}{2}\log 2\pi-\frac{d}{2}\log\tau g(t)^{2}-\frac{d}{2}$
$-\frac{1}{2}\tau g(t)^{2}\mathbb{E}_{p_{D}}\left[\left\|\nabla\log p_{D}(z_{t})\right\|^{2}\right]$
$-\frac{1}{2}\tau g(t)^{2}\mathbb{E}_{p_{D}}\left[\left\|s\left(z_{t},x\right)\right\|^{2}\right]$
$+\mathbb{E}_{p_{D}}\left[\left(z_{t-\tau}-z_{t}+f(z_{t},t)\tau\right)^{\top}s\left(z_{t},x\right)\right],$
where $(i)$ results from the independence between $\bar{u}_{t}$ and $z_{t}$.
By using (9) and the first-order Taylor expansion, we have
$z_{t-\tau}-z_{t}+f(z_{t},t)\tau$ $=$ $z_{t-\tau}-\left(z_{t-\tau}+f(z_{t-\tau},t-\tau)\tau+\sqrt{\tau}g(t-\tau)u_{t-\tau}\right)$
$+f\left(z_{t-\tau}+f(z_{t-\tau},t)\tau+\sqrt{\tau}g(t-\tau)u_{t-\tau},t\right)\tau$
$=$ $-f(z_{t-\tau},t-\tau)\tau-\sqrt{\tau}g(t-\tau)u_{t-\tau}$
$+\tau f(z_{t-\tau},t-\tau)$
$+\tau\frac{\partial f(z_{t-\tau},t-\tau)}{\partial z_{t-\tau}}\left(z_{t}-z_{t-\tau}\right)$
$+\tau^{2}\frac{\partial f(z_{t-\tau},t-\tau)}{\partial t}$
$+\tau\cdot O\left(\left\|z_{t}-z_{t-\tau}\right\|^{2}+\tau^{2}\right)$
$=$ $-\sqrt{\tau}g(t-\tau)u_{t-\tau}$
$+\tau\frac{\partial f(z_{t-\tau},t-\tau)}{\partial z_{t-\tau}}\left(f(z_{t-\tau},t-\tau)\tau+\sqrt{\tau}g(t-\tau)u_{t-\tau}\right)$
$+\tau^{2}\frac{\partial f(z_{t-\tau},t-\tau)}{\partial t}$
$+\tau\cdot O\left(\left\|f(z_{t-\tau},t-\tau)\tau+\sqrt{\tau}g(t-\tau)u_{t-\tau}\right\|^{2}+\tau^{2}\right)$
$=$ $-\sqrt{\tau}g(t-\tau)u_{t-\tau}+O(\tau^{2})$
and
$s\left(z_{t},x,t\right)$ $=$ $s(z_{t-\tau},x,t-\tau)+\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial z_{t-\tau}}\left(z_{t}-z_{t-\tau}\right)+\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial t}\tau$
$=$ $s(z_{t-\tau},x,t-\tau)$
$+\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial z_{t-\tau}}\left(f(z_{t-\tau},t-\tau)\tau+\sqrt{\tau}g(t-\tau)u_{t-\tau}\right)$
$+\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial t}\tau$
$=$ $s(z_{t-\tau},x,t-\tau)+\sqrt{\tau}\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial z_{t-\tau}}g(t-\tau)u_{t-\tau}+O(\tau).$
Combining the two equations above, we get
$\mathbb{E}_{p_{D}}\left[\left(z_{t-\tau}-z_{t}+f(z_{t},t)\tau\right)^{\top}s\left(z_{t},x,t\right)\right]$ $=$ $\mathbb{E}_{p_{D}}\left[-\sqrt{\tau}g(t-\tau)u_{t-\tau}^{\top}s(z_{t-\tau},x,t-\tau)\right]$ (14)
$-\tau g(t-\tau)^{2}\mathbb{E}_{p_{D}}\left[u_{t-\tau}^{\top}\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial z_{t-\tau}}u_{t-\tau}\right]$
$+o(\tau)$
=$\overset{(ii)}{=}$ $-\tau g(t-\tau)^{2}\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial z_{t-\tau}}\right)\right]$
$+o(\tau),$
where ($ii$) follows from the fact that $u_{t-\tau}$ is independent of $z_{t-\tau}$ and $x$ in the decoding process. Taking into account (14),(12) and (13), we arrive at the following results:
$\Delta_{t}$$\triangleq$ $\mathbb{E}_{p_{D}}\left[\log p_{D}(z_{t-\tau}|z_{t})\right]-\mathbb{E}_{p_{D}}\left[\log p_{E}(z_{t-\tau}|z_{t},x)\right]$
$=$ $\frac{1}{2}\tau g(t)^{2}\mathbb{E}_{p_{D}}\left[\left\|\nabla\log p_{D}(z_{t})\right\|^{2}\right]$
$+\frac{1}{2}\tau g(t)^{2}\mathbb{E}_{p_{D}}\left[\left\|s\left(z_{t},x,t\right)\right\|^{2}\right]$
$+\tau g(t-\tau)^{2}\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t-\tau},x,t-\tau)}{\partial z_{t-\tau}}\right)\right]+o(\tau).$
Let $\tau=T/L$, where $L$ is a large number, we have
$\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x,z_{0},z_{\tau},\ldots,z_{T})}{p_{E}(x,z_{0},z_{\tau},\ldots,z_{T})}\right]$ $=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0})p_{D}(z_{T})\prod_{l=1}^{L}p_{D}\left(z_{(l-1)\tau}|z_{l\tau}\right)}{\pi(x)p_{E}(z_{T})\prod_{l=1}^{L}p_{E}\left(z_{(l-1)\tau}|z_{l\tau},x\right)}\right]$
$=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0})}{Z^{-1}\exp\left(-U(x)\right)}\right]+\sum_{l=1}^{L}\Delta_{l\tau}$
$=$ $\mathbb{E}_{p_{D}}\left[\log p_{D}(x|z_{0})+U(x)\right]+\log Z+\sum_{l=1}^{L}\Delta_{l\tau}$
and
$\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x,z_{[\cdot]})}{p_{E}(x,z_{[\cdot]})}\right]$ $=$ $\lim_{L\to\infty}\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x,z_{0},z_{\tau},\ldots,z_{T})}{p_{E}(x,z_{0},z_{\tau},\ldots,z_{T})}\right]$
$=$ $\mathbb{E}_{p_{D}}\left[\log p_{D}(x|z_{0})+U(x)\right]$
$+\int\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\left[\left\|s\left(z_{t},x,t\right)\right\|^{2}+2\mathrm{tr}\left(\frac{\partial s(z_{t},x,t)}{\partial z_{t}}\right)+\left\|\nabla\log p_{D}(z_{t})\right\|^{2}\right]\mathrm{d}t$
$+\log Z.$
Considering integration by parts, $D_{KL}(p_{D}||p_{E})$ in Eq.(6) has the following equivalent expression
@@
$=$ $-\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)|z_{0}\right]$
$\Rightarrow\mathbb{E}_{p_{D}}\left[s\left(z_{t},x\right)^{\top}\nabla_{z_{t}}\log p_{D}(z_{t}|z_{0})\right]$ $=$ $-\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)\right]$
Substituting the above equation to Eq.(6), we have
$D_{KL}\left(p_{D}(x,z_{[\cdot]})||p_{E}(x,z_{[\cdot]})\right)$ $=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0};\phi)}{\pi(x)}\right]$

### v3

@@ 4 Experiments
$T=2.0$ $260.24\pm 0.13$ $210.17\pm 0.43$ $\mathbf{270.32\pm 0.18}$
$T=2.1$ $250.57\pm 0.14$ $208.48\pm 0.41$ $\mathbf{260.34\pm 0.19}$
$T=2.2$ $239.43\pm 0.16$ $210.57\pm 0.39$ $\mathbf{252.11\pm 0.17}$
$T=2.3$ $231.25\pm 0.15$ $214.64\pm 0.37$ $\mathbf{233.46\pm 0.17}$
$T=2.4$ $\mathbf{225.69\pm 0.17}$ $12.43\pm 0.37 22$$25.02\pm 0.15$
[…]
$T=2.6$ $\mathbf{216.69\pm 0.18}$ $181.83\pm 0.40$ $214.78\pm 0.14$
$T=2.7$ $212.42\pm 0.18$ $189.16\pm 0.36$ $\mathbf{212.57\pm 0.14}$
Ising model Finally, we verify the performance of EDG on the 2-dimensional Ising model [ising], a mathematical model of ferromagnetism in statistical mechanics. To ensure the continuity of physical variables, we employ a continuous relaxations trick [isingrelax] to transform discrete variables into continuous auxiliary variables with the target distribution:
$\pi(x)=\exp\left(-\frac{1}{2}x^{T}\left(K(T)+\alpha I\right)^{-1}x\right)\times\prod_{i=1}^{N}\cosh\left(x_{i}\right),$
where $K$ is an $N\times N$ symmetric matrix depending on the temperature $T$, $\alpha$ is a constant guaranteeing $K+\alpha I$ to be positive. For the corresponding discrete Ising variables $\boldsymbol{s}=\{1,-1\}^{\otimes N}$, one can directly obtain discrete samples according to $\pi(\boldsymbol{s}|\boldsymbol{x})=\prod_{i}\left(1+e^{-2s_{i}x_{i}}\right)^{-1}$. When there is no external magnetic field and each spin can only interact with its neighboring spins, $K$ is defined as $\sum_{<ij>}s_{i}s_{j}/T$, with the nears neighboring sum <ij><ij>. Consequently, the normalizing constant of the continuous relaxations system is given by $\log Z=\log Z_{\text{Ising }}+\frac{1}{2}\ln\operatorname{det}(K+\alpha I)-\frac{N}{2}[\ln(2/\pi)-\alpha]$ [ising]. Additionally, using the method described in Sec. 3.3, we provide the estimates of $\log Z_{\text{Ising}}$ at different temperatures for samples generated by NeuralRG, PIS and EDG. Since these are lower bound estimates, larger values indicate more accurate results. As seen in Table 4, EDG provides the most accurate estimates of $\log Z$ across most temperature ranges. The generated states at different temperatures are displayed in Fig. 4.
@@ 5 Conclusion
Our work introduces the EDG as an innovative and effective sampling approach by combining principles from VI and diffusion based methods. Drawing inspiration from VAEs, EDG excels in efficiently generating samples from intricate Boltzmann distributions. Leveraging the expressive power of the diffusion model, our method accurately estimates the KL divergence without the need for numerical solutions to ordinary or stochastic differential equations. Empirical experiments validate the superior sampling performance of EDG.
In consideration of its powerful generation ability and unrestrained network design theoretically, there is still room for further exploration. We can design specific network architectures for different tasks. Regarding many-body particle systems, we plan to leverage equivariant graph neural networks (EGNN) [kohler2020equivariant, klein2023equivariant] to design all trainable components within the GHD module of the decoder to ensure the invariance. Moreover, we aim to construct equivariant score function models within the encoding process with respect to both $x$ and $z_{t}$, which is expected to guarantee the invariance of the weight function. Furthermore, theoretical convergence guarantees can be strengthened by investigating error bounds and convergence rates under finite training time and limited model capacity.
In summary, our future work will extend the application of EDG to construct generative models for large-scale physical and chemical systems, such as proteins [protein-1, protein-2], while simultaneously refining the theoretical foundations of our model.
@@ Acknowledgements
The first and third authors are supported by the NSF of China (under grant number 12171367). The second author is supported by the NSF of China (under grant numbers 92270115, 12071301) and the Shanghai Municipal Science and Technology Commission (No. 20JC1412500), and Henan Academy of Sciences. The last author was supported by the NSF of China (No. 12288201), and the Research Startup Fund of Henan Academy of Sciences (No. 232019024).
Appendix A Proof of Eq. (1)
For completeness, here we provide a proof of Eq. (1), which follows a similar derivation as the loss function of the standard VAE [kingma2019introduction]. For simplicity of notation, we omit the explicit dependence on the distribution parameters $\phi$ and $\theta$.
The KL divergence between $p_{D}(x,z_{0})$ and $p_{E}(x,z_{0})$ can be written as
$D_{\mathrm{KL}}\left(p_{D}(x,z_{0})||p_{E}(x,z_{0})\right)$ $=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x,z_{0})}{p_{E}(x,z_{0})}\right]$
$=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x)}{\pi(x)}\right]+\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(z_{0}|x)}{p_{E}(z_{0}|x)}\right]$
$=$ $D_{\mathrm{KL}}\left(p_{D}(x)||\pi(x)\right)$
$+D_{\mathrm{KL}}\left(p_{D}(z_{0}|x)||p_{E}(z_{0}|x)\right),$
where
$D_{\mathrm{KL}}\left(p_{D}(z_{0}|x)\,\|\,p_{E}(z_{0}|x)\right)\triangleq\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(z_{0}|x)}{p_{E}(z_{0}|x)}\right]$
is the expected KL divergence between the conditional distributions $p_{D}(z_{0}|x)$ and $p_{E}(z_{0}|x)$ under $x\sim p_{D}(x)$, and is therefore non-negative.
Consequently, we obtain
$D_{\mathrm{KL}}\left(p_{D}(x)\,\|\,\pi(x)\right)$ $\leq$ $D_{\mathrm{KL}}\left(p_{D}(x,z_{0})\,\|\,p_{E}(x,z_{0})\right)$
$=$ $D_{\mathrm{KL}}\left(p_{D}(z_{0})p_{D}(x|z_{0})\,\|\,\pi(x)p_{E}(z_{0}|x)\right),$
where the inequality becomes an equality if $p_{D}(z_{0}|x)=p_{E}(z_{0}|x)$ for all $x,z_{0}$.
Appendix B Analysis of the independence between $z_{T}$ and $x$
In the decoding process, if $z_{T}$ is independent of $z_{0}$ with $p_{D}(z_{T}|z_{0})=p_{D}(z_{T})$, we have
$p_{D}(x,z_{T})$ $=$ $\int p_{D}(x,z_{0},z_{T})\mathrm{d}z_{0}$
$=$ $\int p_{D}(x,z_{0})p_{D}(z_{T}|z_{0})\mathrm{d}z_{0}$
$=$ $p_{D}(z_{T})\int p_{D}(x,z_{0})\mathrm{d}z_{0}$
$=$ $p_{D}(x)p_{D}(z_{T}),$
which implies that $x$ is also independent of $z_{T}$.
Appendix C Proof of Theorem 1
Part 1
Eq. (6) can be proved in a manner similar to the proof of Eq. (1) (see A). Thus, here we focus on proving that the equality in Eq. (6) holds under the conditions that $z_{0}$ and $z_{T}$ are independent, and $s(z,x,t;\theta)\equiv\nabla_{z_{t}}\log p_{D}(z_{t}|x)$.
Under these assumptions, $x$ is also independent of $z_{T}$ in the decoding process (see B). Consequently, the conditional distribution of $z_{[\cdot]}$ given $x$ can be described by the reverse-time process defined by Eq. (3). Moreover, it shares the same initial distribution $p_{D}(z_{T}|x)=p_{D}(z_{T})=p_{E}(z_{T})$ as the reverse-time process defined by Eq. (4) in the encoding process.
Furthermore, because $s(z,x,t;\theta)\equiv\nabla_{z_{t}}\log p_{D}(z_{t}|x)$, the drift and diffusion terms of the two reverse-time processes exactly match. Therefore, we have $p_{D}(z_{[\cdot]}|x)=p_{E}(z_{[\cdot]}|x)$, and
$D_{\mathrm{KL}}\left(p_{D}(z_{[\cdot]},x)\,\|\,p_{E}(z_{[\cdot]},x)\right)$ $=$ $D_{\mathrm{KL}}\left(p_{D}(x)\,\|\,\pi(x)\right)$
$+D_{\mathrm{KL}}\left(p_{D}(z_{[\cdot]}|x)\,\|\,p_{E}(z_{[\cdot]}|x)\right)$
$=$ $D_{\mathrm{KL}}\left(p_{D}(x)\,\|\,\pi(x)\right).$
Part 2
We now prove Eq. (6). In the decoding process, the distribution of $z_{[\cdot]}$ can be described by the following reverse-time process:
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}})\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}\tilde{W}_{\tilde{t}},\ z_{T}\sim p_{D}(z_{T}).$
Combining this with Eq. (4) and applying Girsanov theorem [trick-1], we obtain
@@ Appendix C Proof of Theorem 1
$\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(z_{[\cdot]})}{p_{E}(z_{[\cdot]}|x)}\right]$ $=$ $\mathbb{E}_{p_{D}}\left[\int_{0}^{T}g(t)\left(\nabla_{z_{t}}\log p_{D}(z_{t})-s(z_{t},x,t;\theta)\right)\mathrm{d}\tilde{W}_{t}\right]$
$+\frac{1}{2}\mathbb{E}_{p_{D}}\left[\int_{0}^{T}g(t)^{2}\left\|\nabla_{z_{t}}\log p_{D}(z_{t})-s(z_{t},x,t;\theta)\right\|^{2}\mathrm{d}t\right].$
Note that the first term on the right-hand side equals zero due to the martingale property of Itô integrals. Thus, we have
$D_{\mathrm{KL}}\left(p_{D}(z_{[\cdot]},x)\,\|\,p_{E}(z_{[\cdot]},x)\right)$ $=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0})}{\pi(x)}\right]+\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(z_{[\cdot]})}{p_{E}(z_{[\cdot]}|x)}\right]$ (15)
$=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0})}{\pi(x)}\right]$
$+\frac{1}{2}\mathbb{E}_{p_{D}}\left[\int_{0}^{T}g(t)^{2}\left\|\nabla_{z_{t}}\log p_{D}(z_{t})-s(z_{t},x,t;\theta)\right\|^{2}\mathrm{d}t\right]$
$=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0})}{\pi(x)}\right]$
$+\frac{1}{2}\mathbb{E}_{p_{D}}\left[\int_{0}^{T}g(t)^{2}\left(\left\|\nabla_{z_{t}}\log p_{D}(z_{t})\right\|^{2}+\left\|s(z_{t},x,t;\theta)\right\|^{2}\right)\mathrm{d}t\right]$
$+\mathbb{E}_{p_{D}}\left[\int_{0}^{T}g(t)^{2}\left(s(z_{t},x,t;\theta)^{\top}\nabla_{z_{t}}\log p_{D}(z_{t})\right)\mathrm{d}t\right].$
Moreover, we consider the following integration by parts:
@@
$=$ $-\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)|z_{0}\right],$
which yields
$\mathbb{E}_{p_{D}}\left[s\left(z_{t},x\right)^{\top}\nabla_{z_{t}}\log p_{D}(z_{t}|z_{0})\right]$ $=$ $-\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)\right].$ (16)
By substituting Eq. (16) into Eq. (15), we finally obtain Eq. (6).
It is worth noting that we could also directly design a loss function based on Eq. (15). Such a loss is theoretically equivalent to the one used in this work, but may suffer from large variance in $\nabla_{z_{t}}\log p_{D}(z_{t})$ when $t$ is small.
Appendix D Proof of equivalence between Eq. (7) and Eq. (8)
First, we prove the Hutchinson estimator in Eq. (8). In Eq. (7), one frequently needs the divergence $\nabla\cdot s(z_{t},x,t;\theta)=\mathrm{tr}(\partial_{z_{t}}s)$ of a vector field $s:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Deep learning frameworks (PyTorch, TensorFlow) do not natively provide an efficient way to form an entire $d\times d$ Jacobian for large $d$, while Hutchinson estimator replace $\mathrm{tr}(\cdot)$ with the unbiased Monte Carlo estimate without full Jacobians.
Considering a random variable $p(\epsilon)\sim\mathrm{Rademacher}^{d}$ with $\mathbb{E}[\epsilon]=0$ and $\mathrm{Cov}(\epsilon)=I$, we can get a stochastic estimation of the trace according to [hutchinson1989stochastic]
$\mathrm{tr}(\partial_{z_{t}}s)$ $=$ $\mathbb{E}_{p(\epsilon)}[\epsilon^{\top}\partial_{z_{t}}s\ \epsilon].$
By the rules of matrix differential calculus,
$\mathrm{d}\bigl(\epsilon^{\top}s(z_{t},x,t;\theta)\bigr)=\epsilon^{\top}\,\mathrm{d}s(z_{t},x,t;\theta)=\epsilon^{\top}\,\bigl(\partial_{z_{t}}s(z_{t},x,t;\theta)\bigr)\,\mathrm{d}z_{t}.$
Then, it enables the use of automatic differentiation for calculation,
$\epsilon^{\top}\,\bigl(\partial_{z_{t}}s(z_{t},x,t;\theta)\bigr)\,=\frac{\partial\bigl[\epsilon^{\top}s(z_{t},x,t;\theta)\bigr]}{\partial z_{t}}.$
Therefore, we have
$\nabla\cdot s(z_{t},x,t;\theta)$ $=$ $\mathbb{E}_{p(\epsilon)}[\frac{\partial\bigl[\epsilon^{\top}s(z_{t},x,t;\theta)\bigr]}{\partial z_{t}}\ \epsilon],$
where $p(\epsilon)\sim\mathrm{Rademacher}^{d}$ denotes a random vector in $\mathbb{R}^{d}$ with i.i.d. Rademacher entries, satisfying $\mathbb{E}[\epsilon]=0$ and $\mathrm{Cov}(\epsilon)=I$.
Then, we prove the expectation of a uniform time distribution $t$ over $[0,T]$ in Eq. (8). When $t\sim\mathcal{U}[0,T]$, its probability density function is
$p_{\mathcal{U}[0,T]}(t)=\begin{cases}\frac{1}{T},&0\leq t\leq T\\ 0,&\text{ otherwise }\end{cases}\ .$
Then we simplify the integration in Eq. (7)
$\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}[\cdot]\mathrm{d}t$ $=$ $\int_{0}^{T}\frac{1}{T}\frac{Tg(t)^{2}}{2}\mathbb{E}_{p_{D}}[\cdot]\mathrm{d}t$
$=$ $\int_{-\infty}^{+\infty}p_{\mathcal{U}[0,T]}(t)\frac{Tg(t)^{2}}{2}\mathbb{E}_{p_{D}}[\cdot]\mathrm{d}t$
$=$ $\mathbb{E}_{t\sim\mathcal{U}[0,T]}\left[\frac{Tg(t)^{2}}{2}\mathbb{E}_{p_{D}}[\cdot]\right].$
Combining two parts above, we can achieve